OpenStudy (jakeybabs):

-49=6x-13-4x what is x?

OpenStudy (anonymous):

49 = 6x-13-4x 49= 2x-13 62= 2x x= 31 therefor the answer is: x = 31

OpenStudy (anonymous):

Rearrange the eqation by puttin the x on one side an the constants on the other.

OpenStudy (jakeybabs):

its negative 49 though

OpenStudy (anonymous):

Your right, so what are you having troubles with here?

OpenStudy (anonymous):

-49 = 6x-13-4x -49= 2x-13 -36= 2x x= -18 therefor x = -18 very sorry, i didn't notice the negative :o

OpenStudy (jakeybabs):

lol no prob, thank you so much for explaining the steps :)